OpenStudy (anonymous):

What is the equation of the line perpendicular to x+5y=-2 passing through (-3,-4) in standard form?

OpenStudy (hoblos):

slope*(-1/5)=-1 slope=5 y+4=5(x+3) y=5x+11

OpenStudy (campbell_st):

find the gradient of your line.... then the gradient of the perpendicular is -1/m then use point gradient to find the equation..

OpenStudy (anonymous):

that is in slope-intercept how do I put y=5x+11 in standard ?

OpenStudy (hoblos):

5x-y=-11

OpenStudy (campbell_st):

so everything on the left hand side.... y - 5x -11 = 0

OpenStudy (anonymous):

moving the 11 to the left hand side would make it positive right?

OpenStudy (anonymous):

i think I will go with 5x-y=-11, thanks.

OpenStudy (hoblos):

yup thats it my pleasure XD